OpenStudy (anonymous):

A bag contains 5 red marbles, 6 white marbles and 5 blue marbles. Find P(red or blue)

OpenStudy (anonymous):

p is 5

OpenStudy (anonymous):

thank you

OpenStudy (kropot72):

The events 'draw red' and 'draw blue' are mutually exclusive. Therefore the required probability is the sum of the individual probabilities. \[P(red\ or\ blue)=\frac{5}{16}+\frac{5}{16}=you\ can\ calculate\]

OpenStudy (anonymous):

P(red or blue) = P(red) + P(blue) = 5/16 + 5/16

OpenStudy (anonymous):

Then to find the experimental probability if needed simplify the sum from \[\frac{ 10 }{ 16? }\] to get your experimental probability.

OpenStudy (anonymous):

Forget that question mark it got in the way.

OpenStudy (anonymous):

Do you mean in decimal?

OpenStudy (anonymous):

In fraction

OpenStudy (anonymous):

Like the experimental probability of 8\[\frac{ 80 }{ 100 }\] is \[\frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

Forget that 8 also after of

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Was a mistake

OpenStudy (anonymous):

that means you need to simplify 10/16 = (2*5)/(2*8) = 5/8

OpenStudy (anonymous):

Just a sec

OpenStudy (anonymous):

Yes correct

OpenStudy (anonymous):

Gave you medal